# arXiv:2602.22118v2[cs.RO] 27 Feb 2026

### 27 Feb 2026 v2

@@ Results
The robot’s main computer is an UP Xtreme i12, which runs a 1 kHz control loop responsible for reading from and writing to the connected devices (motor drivers, IMUs, encoders, etc.), as shown in Fig. S3. By having three 3-DoF acceleration sensors on each rigid link, we are able to measure the instantaneous linear acceleration, angular acceleration, and angular velocity of each link. (Although this could have been achieved with two IMUs and four accelerometers, reuse of the IMU saved us development time and complexity). This allows the estimation of front and rear ground contact state via a two-body lumped dynamics model that estimates tire contact forces and classifies probable wheel contacts.
@@ Discussion
In order to control UMV, an estimate of all states and their first time derivatives is needed. This complete state is comprised of the 6-DoF floating base state and the six serial joint states. Because the front wheel’s mass is low, its velocity and acceleration can generally be neglected. Of the six serial joint states $q_{h}$, $\mu$, $\phi$, $q_{s}$, $q_{w}$, $\theta$], the position and velocity of the five that are necessary for dynamics are directly measured (i.e., all but $\theta$). The largest state estimation challenge required to drive UMV to the edge of its capability is estimation of the 6-DoF floating base state and its derivative. This state is comprised of the 3-DoF orientation, parameterized in our optimization as a $\times3$ rotation matrix, and the 3-DoF position of the the Bike link frame. The linear velocity (position derivative) must be estimated, as it cannot be directly observed. The angular velocity is directly measured via rate gyro from an IMU and does not require explicit estimation. As an engineering compromise, and to let us focus on platform design and control, we utilize a motion capture system to aid in state estimation while in the lab. Motion capture estimates are fused with a high rate onboard IMU to produce quality base-state estimates. The robot does support self-contained state-estimation, which has been demonstrated in field operations. Retro-reflective markers on the robot are tracked by an off-the-shelf Optitrack motion capture system to produce position and orientation estimates at 120Hz. In order to provide high rate, low latency estimation to controllers, we use the GTSAM [13] factor graph optimization package to fuse the 1kHz IMU signal and the lower rate Optitrack signal. We adopted a bipartite factor graph architecture [18] where the IMU angular rates and linear accelerations are integrated at 1kHz to provide a smooth low latency pose estimate, and the Optitrack measurements are decimated to 30Hz and used to correct this high frequency estimate. A factor graph tracks position, orientation, linear velocity, and gyro and accelerometer biases. We also use both position and orientation measurements from the motion capture system. Because our system is highly dynamic and sometimes generates significant noise or saturation in the IMU, we find including the orientation estimates in the graph improves the robustness of the estimator. The estimator is updated at 1 kHz and optimized at 30Hz using the latest motion capture measurement and pre-integrated IMU measurements [14]. This system provides high quality, low latency state information suitable for control of a highly dynamic mechanism. In particular, we find that accurate linear velocity estimates enhance the performance of dynamic balancing controllers, such as the wheelie and hopping controllers.
@@ Acknowledgments
The authors would like to thank Abel Gawel, Aleksandrs Ecins, Ananya Agarwal, Jessica Hodgins, Katie Hughes, Luca Buonocore, Marco Zenti, Petr Listov, and Victor Reijgwart for their technical contributions, dedicated support, and insightful discussions throughout the development and testing of the UMV platform.
@@
Author contributions: The Core Contributors developed and implemented the design optimization pipeline, designed and assembled the UMV robot, developed the embedded and training-tool software, trained most of the policies, and oversaw the hardware experiments on robots. They also contributed to the writing of the manuscript. They assisted with engineering tasks to support the project, including deployment code on hardware, data collection and curation, as well as supporting hardware experiments. The Project Leads set the overarching vision by defining the core scientific goals and managed the project.
@@ A Driving, Balancing, and Jumping Bicycle Robot
Benjamin Bokser, Daniel J. Gonzalez, Surya P. N. Singh, Aaron Preston, Alex Bahner, Annika Wollschläger, Arianna Ilvonen, Asa Eckert-Erdheim, Ashwin Khadke, Bilal Hammoud, Dean Molinaro, Fabian Jenelten, Henry Mayne, Howie Choset, Igor Bogoslavskyi, Itic Tinman, James Tigue, Jan Preisig, Kaiyu Zheng, Kenny Sharma, Kim Ang, Laura Lee, Liana Margolese, Nicole Lin, Oscar Frias, Paul Drews, Ravi Boggavarapu, Rick Burnham, Samuel Zapolsky, Sangbae Kim, Scott Biddlestone, Sean Mayorga, Shamel Fahmi, Tyler McCollum, Velin Dimitrov, William Moyne, Yu-Ming Chen, David Perry, Farbod Farshidian, Marco Hutter, Al Rizzi, and Gabe Nelson

### 17 Mar 2026 v3

# System Design of the Ultra Mobility Vehicle: A Driving, Balancing, and Jumping Bicycle Robot

@@ Results
The robot’s main computer is an UP Xtreme i12, which runs a 1 kHz control loop responsible for reading from and writing to the connected devices (motor drivers, IMUs, encoders, etc.), as shown in Fig. S3. By having three 3-DoF acceleration sensors on each rigid link, we are able to measure the instantaneous linear acceleration, angular acceleration, and angular velocity of each link. (Although this could have been achieved with two IMUs and four accelerometers, reuse of the IMU saved us development time and complexity.) This allows the estimation of front and rear ground contact state via a two-body lumped dynamics model that estimates tire contact forces and classifies probable wheel contacts.
@@ Discussion
In order to control UMV, an estimate of all states and their first time derivatives is needed. This complete state is comprised of the 6-DoF floating base state and the six serial joint states. Because the front wheel’s mass is low, its velocity and acceleration can generally be neglected. Of the six serial joint states $q_{h}$, $\mu$, $\phi$, $q_{s}$, $q_{w}$, $\theta$], the position and velocity of the five that are necessary for dynamics are directly measured (i.e., all but $\theta$). The largest state estimation challenge required to drive UMV to the edge of its capability is estimation of the 6-DoF floating base state and its derivative. This state is comprised of the 3-DoF orientation, parameterized in our optimization as a $\times3$ rotation matrix, and the 3-DoF position of the Bike link frame. The linear velocity (position derivative) must be estimated, as it cannot be directly observed. The angular velocity is directly measured via rate gyro from an IMU and does not require explicit estimation. As an engineering compromise, and to let us focus on platform design and control, we utilize a motion capture system to aid in state estimation while in the lab. Motion capture estimates are fused with a high rate onboard IMU to produce quality base-state estimates. The robot does support self-contained state-estimation, which has been demonstrated in field operations. Retro-reflective markers on the robot are tracked by an off-the-shelf Optitrack motion capture system to produce position and orientation estimates at 120Hz. In order to provide high rate, low latency estimation to controllers, we use the GTSAM [13] factor graph optimization package to fuse the 1kHz IMU signal and the lower rate Optitrack signal. We adopted a bipartite factor graph architecture [18] where the IMU angular rates and linear accelerations are integrated at 1kHz to provide a smooth low latency pose estimate, and the Optitrack measurements are decimated to 30Hz and used to correct this high frequency estimate. A factor graph tracks position, orientation, linear velocity, and gyro and accelerometer biases. We also use both position and orientation measurements from the motion capture system. Because our system is highly dynamic and sometimes generates significant noise or saturation in the IMU, we find including the orientation estimates in the graph improves the robustness of the estimator. The estimator is updated at 1 kHz and optimized at 30Hz using the latest motion capture measurement and pre-integrated IMU measurements [14]. This system provides high quality, low latency state information suitable for control of a highly dynamic mechanism. In particular, we find that accurate linear velocity estimates enhance the performance of dynamic balancing controllers, such as the wheelie and hopping controllers.
@@ Acknowledgments
The authors would like to thank Abel Gawel, Aleksandrs Ecins, Ananya Agarwal, Elliott Rouse, Jessica Hodgins, Katie Hughes, Luca Buonocore, Marco Zenti, Petr Listov, and Victor Reijgwart for their technical contributions, dedicated support, and insightful discussions throughout the development and testing of the UMV platform. We also acknowledge the excellent interns who contributed to the UMV project: Abby Speckhals, Alexandre Amice, Antoine Leeman, Ardalan Tajbakhsh, Ashley Su, Catrina Sarno, Chris Evagora, Jeonghwan Kim, Ryan Grummer, and Seungeun Rho.
@@
Author contributions: The Contributors developed and implemented the design optimization pipeline, designed and assembled the UMV robot, developed the embedded and training-tool software, trained most of the policies, and oversaw the hardware experiments on robots. They also contributed to the writing of the manuscript. They assisted with engineering tasks to support the project, including deployment code on hardware, data collection and curation, as well as supporting hardware experiments. The Project Leads set the overarching vision by defining the core scientific goals and managed the project.
@@ A Driving, Balancing, and Jumping Bicycle Robot
Benjamin Bokser, Daniel J. Gonzalez, Aaron Preston, Alex Bahner, Annika Wollschläger, Arianna Ilvonen, Asa Eckert-Erdheim, Ashwin Khadke, Bilal Hammoud, Dean Molinaro, Fabian Jenelten, Henry Mayne, Howie Choset, Igor Bogoslavskyi, Itic Tinman, James Tigue, Jan Preisig, Kaiyu Zheng, Kenny Sharma, Kim Ang, Laura Lee, Liana Margolese, Nicole Lin, Oscar Frias, Paul Drews, Ravi Boggavarapu, Rick Burnham, Samuel Zapolsky, Sangbae Kim, Scott Biddlestone, Sean Mayorga, Shamel Fahmi, Surya P. N. Singh, Tyler McCollum, Velin Dimitrov, William Moyne, Yu-Ming Chen, David Perry, Farbod Farshidian, Marco Hutter, Al Rizzi, and Gabe Nelson